OpenStudy (willa):

Which of the following is a factor of x2 - 7x - 8? A) x + 1 B) x + 8 C) x - 8 D) x - 1 E) x - 7

OpenStudy (amistre64):

c

OpenStudy (amistre64):

a and c really

OpenStudy (willa):

how do you figure it out???

OpenStudy (amistre64):

well; you have to factor it apart; and you have to know what terms are produced how. the last term is multiplied to get -8 and the middle term sums those 2 together -8+1 = -7 -8*1 = -8

OpenStudy (amistre64):

(x-8)(x+1) = x^2 -7x -8 FOIL is the terminology they use to work it back

OpenStudy (willa):

ok thankk you sooooooooooooo much

OpenStudy (amistre64):

youre welcome :)